OpenStudy (anonymous):

find the value of n such that x^2-11x+n is a perfect square trinomial

OpenStudy (anonymous):

So basically, you're looking for two terms that when multiplied by each other (Squared), give us 'n' and when added to each other (same as multiplied by 2), give the value of -11. Let's call this number u; such that when x is multiplied by 2 gives -11 and when squared gives n.\[\bf 2u=-11\]\[\bf u^2=n\]Let's rearrange the first equation and solve for u:\[\bf 2u=-11 \implies u =-\frac{ 11 }{ 2 }\]Let's plug this value you in for u in the second equation and solve for n:\[\bf u^2=n \rightarrow \left( -\frac{ 11 }{ 2 } \right)^2=n \implies n = ?\] Can you now solve for n?

OpenStudy (anonymous):

@mslee0105

geerky42 (geerky42):

@genius12

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

In the last sentence where it says: "...such that when x is multiplied by 2 gives...", replace the 'x' with 'u'.